# Electron-phonon-dominated charge-density-wave fluctuations in TiSe2 accessed by ultrafast nonequilibrium dynamics

### v1

@@ III Ultrafast Melting and Recovery of RT-CDW Fluctuations
Another recent ultrafast low-energy electron diffuse scattering study investigated the melting process of charge order in TiSe2 and estimated roughly 30:70 ratio of excitonic versus electron-phonon contributions to the total lattice distortion in the low-temperature CDW phase Kurtz et al. (2024). These results imply that although short-range CDW fluctuations are a dynamical phenomenon fluctuating in time and space, the microscopic mechanisms that stabilize these ordered quantum states and govern their ultrafast light-induced dynamics are the same as in the conventional (low-temperature) CDW phases. These short-range CDW fluctuations and their coexistence with some normal (non-CDW) domains can explain the low contrast of the coherent oscillations associated with the CDW amplitude phonon mode. Indeed, in low-temperature standard CDW phases, high-contrast modulations were measured by tr-ARPES Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021).
IV Theoretical Analysis of Dynamical Electron-Phonon Interactions
@@
To further disentangle the influence of the electron-phonon coupling on the bare electronic structure above the transition temperature, we turn our attention to the theoretical framework that includes electron self-energy corrections necessary for the inclusion of electron-phonon interaction. Notably, no term responsible for the electron-hole coupling was used here. Yet, the spectral function strikingly matches the experimentally measured features that we attribute to the CDW-fluctuating phase. In Fig. 4, we show the impact of dynamical electron-phonon scattering on the electron and phonon properties of 1T-TiSe2 when the system is just above the structural instability . Panels (a)-(d) show the electron spectral functions along the -M high symmetry direction and CECs (at Fermi energy) in the first Brillouin zone. The results for a temperature above but close to the CDW instability are depicted in Fig. 4 (a) and (b), while the other two panels are for . The electron-phonon interaction is included into the spectral function by using the dynamical Fan-Migdal electron self-energy Giustino (2017); Lee et al. (2023) based on inputs derived within the density functional perturbation theory Baroni et al. (2001); Giannozzi et al. (2017). Even though we use the original unit cell with no CDW periodic lattice distortions, we observe strong renormalizations of electron bands for very similar to the characteristic modifications observed in the CDW state, i.e., (i) finite spectral weight below the Ti 3d conduction band at the M point, coming from replicas of Se 4p states at the point, (ii) the flower-shaped contour at the point close to the Fermi level coming from the replicas of the three non-equivalent Ti 3d M pockets, and (iii) the loss of spectral weight at the particular momentum of the M pocket close to the Fermi level [see Fig. 1 (c)]. These signatures come from the electron-phonon fluctuations and very much resemble the CDW fluctuations observed in our ARPES data at RT. With the rise of the temperature, these electron-phonon fluctuations are reduced [see Fig. 4(e) and Fig. S2 in SM], and the electronic band structure becomes closer to the bare DFT result [red curves in Fig. 4(c)]. We observe that these features are related to the soft acoustic phonon mode at the M point [see Fig.4(f)] and are strongest when the energy of this mode is close to zero, while the electron-phonon coupling is the strongest (see Fig. S2 in SM). When the temperature is elevated, the M-point acoustic phonon hardens, which quenches the electron-phonon scatterings between Se 4p and Ti 3d states. This further supports the correlation discussed above between the melting and recovery dynamics of the CDW fluctuating Se 4p band and hardening of the M phonon mode Otto et al. (2021). Similar microscopic explanation, which includes an active soft bosonic mode, was used many years ago to explain the fluctuating (or so-called pseudogap) region above the CDW phase Sadovskii (2001) in the quasi-1D Peierls systems Lee et al. (1973); McKenzie (1995), as well as in the high- cuprates Pines (1997), and more recently in metallic 2H-NbSe2 and 2H-TaSe2 Kuchinskii et al. (2012). Further, spectral weight redistribution around the Fermi level due to electron-phonon-related fluctuations for was once discussed for 1T-TiSe2 Yoshiyama et al. (1986), but it was never thoroughly appreciated and later connected to the replicas and other photoemission signatures in 1T-TiSe2. Our microscopic interpretation is also in line with the observation that the CDW fluctuations reach a maximum at , as observed with X-ray thermal diffuse scattering Holt et al. (2001). The temperature dependence of the phonon scattering lifetime characterized by hardening of the soft phonon mode with increasing temperature and a concomitant reduction of the electron scattering rates (i.e., ), consistent with transport measurements Velebit et al. (2016), is shown in Figs. S2 and S3. It shows the proportionality between the larger CDW gap and the longer lifetime of the amplitude mode. This is most likely caused by less quasiparticle damping and suggests that even in the short-range ordered CDW regions, there is a finite CDW gap that suppresses the damping of the coherent (and re-hardened) amplitude mode oscillation. Notably, the figures also show that artificial introduction of small periodic lattice distortions leads to an increase in the amplitude mode lifetime and better lifetime agreement with the experiment. However, the dynamical electron-phonon scattering alone redistributes the density of states enough to account for a pseudogap formation, as shown in Fig. S3 in SM. We therefore conclude that while static CDW gaps vanish above the transition temperature, the dynamic pseudogap persists. This affects both the electronic scattering and phonon suppression rates.
@@
Figure 4(f) additionally shows the phonon spectral function of 1T-TiSe2 in the fluctuation regime with included phonon-electron coupling, where part of the M phonon spectral weight extends up to negative frequencies. This indicates that in the fluctuation region where , the atomic motions fluctuate between normal and CDW states. Note that equilibrium electron diffraction measurements Cheng et al. (2022) as well as X-ray diffuse scattering Woo et al. (1976b); Guo et al. (2025) observed both L and M diffusive peaks at RT, suggesting the presence of the fluctuating periodic lattice distortions above . The system could, therefore, support the excitation of the CDW amplitude mode, as it is observed experimentally following optical photoexcitation [Fig. 3(f)]. These theoretical results demonstrate that the features (i)-(iii) as well as the CDW amplitude mode observed in the CDW fluctuation regime could be explained by the dynamical electron-phonon effects (also supported by our scanning tunneling microscopy measurements - see Fig. S4 in SM), which opposes the common belief that the fluctuation regime is dominated by the electron-hole correlations Monney et al. (2012, 2015). The latter mechanism could still be present in the formation process of the CDW fluctuations in 1T-TiSe2, but more likely as a higher-order correcting effect. This is further corroborated by the fact that the equivalent electron-related collective soft mode (e.g., plasmon or exciton), which could produce the above features in scattering with Ti 3d or Se 4p states, is not observed in the normal phase above , neither in experiments Lin et al. (2022); Kogar et al. (2017) nor theoretical studies Lian et al. (2019).
V Conclusions
@@
In conclusion, our study provides direct evidence of charge density wave (CDW) fluctuations in 1T-TiSe2 at room temperature using time-resolved extreme ultraviolet (XUV) momentum microscopy. We observe the persistence of the backfolded Se 4p band well above , demonstrating that CDW fluctuations remain detectable even at elevated temperatures. The ultrafast melting and recovery of these fluctuations, featuring the presence of coherent CDW amplitude phonon modes, highlight the importance of electron-phonon interactions in stabilizing the CDW phase.
@@
These results not only refine the understanding of CDW physics in 1T-TiSe2 but also have broader implications for the study of fluctuating phases in other transition metal dichalcogenides and strongly correlated materials. Given the emerging connection between CDW fluctuations and unconventional superconductivity, our findings may provide valuable insights into the interplay between these two phenomena. Future studies leveraging advanced ultrafast techniques could further elucidate the microscopic mechanisms governing the fluctuating CDW phase, potentially unveiling new pathways for manipulating correlated electron systems in quantum materials.
@@ Methods
Experimental Methods
The experimental apparatus is centered around a custom-designed, polarization-tunable ultrafast XUV beamline Comby et al. (2022), interfaced with a Momentum Microscope end-station Medjanik et al. (2017). This beamline is driven by a commercial high-repetition-rate Yb fiber laser (1030 nm, 135 fs FWHM, 50 W) from Amplitude Laser Group, operating at 166 kHz. In the probe branch, a segment of the laser output is frequency-doubled using a BBO crystal, yielding 515 nm pulses (5 W) that are tightly focused into a thin, high-density argon gas jet configured in an annular geometry to facilitate high-order harmonic generation (HHG). Spatial filtering through pinholes effectively separates the intense 515 nm fundamental from the emerging XUV beamlet. To extract the 9th harmonic (21.6 eV), a combination of SiC/Mg multilayer mirrors (NTTAT) and a 200 nm Sn metallic filter is used for spectral selection. On the pump side, a small portion of the original 1030 nm laser pulse (135 fs FWHM) is directed as the IR excitation. The IR pump and XUV probe pulses are recombined collinearly via a drilled mirror and focused onto the target. Bulk iodine vapor transport-grown - single crystals Watson et al. (2019) are freshly cleaved and positioned on a motorized 6-axis hexapod for high-precision alignment within the main vacuum chamber, which maintains an ultra-high base pressure of 210-10 mbar. Photoemission data acquisition is carried out with a custom-designed time-of-flight momentum microscope featuring an advanced front lens capable of multiple operating modes (GST mbH) Tkach et al. (2024b). This detection system enables simultaneous mapping of the entire surface Brillouin zone over a broad binding energy range without necessitating sample orientation adjustments Medjanik et al. (2017). Measurements in this study were conducted at room temperature, where the total energy resolution reaches . This resolution, , results from the convolution of the Fermi-Dirac distribution width at a given temperature , the bandwidth of the ionizing radiation , and the time-of-flight detector resolution . The thermal broadening of the Fermi-Dirac distribution is estimated as , corresponding to approximately at room temperature. For data post-processing, an open-source workflow Xian et al. (2020, 2019) is employed to efficiently convert raw single-event-based datasets into calibrated, binned data hypervolumes, incorporating appropriate axis calibration. Further details about the experimental setup are provided in Ref. Fragkos et al. (2025).
@@
The native defects observed with STM on the sample perfectly match those observed in Arpaia et al. (2019), with an estimated defect density of . Only very few defects of type D associated to intercalated Ti atoms could be observed (density ), validating the low Ti doping of the sample. No signs of CDW modulation have been observed with STM even close to [Figs. S4(a) and S4(b) in SM]. Indeed, the FFT-amplitude plot [Fig. S4(c)] only shows spots corresponding to the unit cell charge modulation.
Computational Methods
Ground state calculations were performed within the framework of the DFT using the Quantum Espresso package Giannozzi et al. (2017). We use optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) with the PBE exchange-correlation functional Perdew et al. (1996) from the PseudoDojo library van Setten et al. (2018). The plane wave energy cutoff was fixed to Ry. To reduce the interaction between the periodic images we modeled the 2D layers in periodic cells with Å vacuum in the direction normal to the layer. The convergence criterion for energy is set to 10-8 eV and the atomic positions are relaxed until the Hellmann-Feynman forces are less than 10-4 eV/Å. A lattice parameter of is obtained with the PBE functional. To obtain the ground state charge density the 2D Brillouin zone is sampled with k-point mesh. The Fermi-Dirac smearing functions were used with the electronic temperatures ranging from 100 K to 1600 K.

### v3

@@ Discussion
Another recent ultrafast low-energy electron diffuse scattering study investigated the melting process of charge order in TiSe2 and estimated roughly 30:70 ratio of excitonic versus electron-phonon contributions to the total lattice distortion in the low-temperature CDW phase Kurtz et al. (2024). These findings highlight that both excitonic effects and electron-phonon coupling play significant roles in stabilizing the CDW order. While short-range CDW fluctuations are a dynamical phenomenon—fluctuating over both time and space—the mechanisms that govern their stabilization and ultrafast dynamics are similar to those in conventional low-temperature CDW phases. These short-range CDW fluctuations and their coexistence with some normal (non-CDW) domains can explain the low contrast of the coherent oscillations associated with the CDW amplitude phonon mode. Indeed, in low-temperature standard CDW phases, high-contrast modulations were measured by tr-ARPES Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021).
@@
Theoretical Analysis of Dynamical Electron-Phonon Interactions. To further disentangle the influence of the electron-phonon coupling on the bare electronic structure above the transition temperature, we turn our attention to the theoretical framework that includes electron self-energy corrections necessary for the inclusion of electron-phonon interaction. Notably, no term responsible for the electron-hole coupling was used here. Yet, the spectral function strikingly matches the experimentally measured features that we attribute to the CDW-fluctuating phase. In Fig. 4, we show the impact of dynamical electron-phonon scattering on the electron and phonon properties of 1T-TiSe2 when the system is just above the structural instability . Temperatures used in the theoretical simulations are scaled with respect to the value from density functional perturbation theory (DFPT), which is 1105 K. The DFPT is the electronic temperature at which the acoustic mode at the M point becomes soft in the harmonic Born-Oppenheimer approximation. Panels (a)-(d) show the electron spectral functions along the -M high symmetry direction and CECs (at Fermi energy) in the first Brillouin zone. The results for a temperature above but close to the CDW instability are depicted in Fig. 4 (a) and (b), while the other two panels are for , when the CDW fluctuating replica are gone and system enters into the normal phase. The electron-phonon interaction is included into the spectral function by using the dynamical Fan-Migdal electron self-energy Giustino (2017); Lee et al. (2023) based on inputs derived within the density functional perturbation theory Baroni et al. (2001); Giannozzi et al. (2017). Even though we use the original unit cell with no CDW periodic lattice distortions, we observe strong renormalizations of electron bands for very similar to the characteristic modifications observed in the CDW state, i.e., (i) finite spectral weight below the Ti 3d conduction band at the M point, coming from replicas of Se 4p states at the point, (ii) the flower-shaped contour at the point close to the Fermi level coming from the replicas of the three non-equivalent Ti 3d M pockets, and (iii) the loss of spectral weight at the particular momentum of the M pocket close to the Fermi level [see Fig. 1 (c)]. These signatures come from the electron-phonon fluctuations and very much resemble the CDW fluctuations observed in our ARPES data at RT. With the rise of the temperature, these electron-phonon fluctuations are reduced [see Fig. 4(e) and Supplementary Figure 2 in SI], and the electronic band structure becomes closer to the bare DFT result [red curves in Fig. 4(c)]. Note that the bimodal intensity profile in our simulations disappears continuously, while it is completely vanished above =1600 K, which is 500 K above the theoretical . We observe that these features are related to the soft acoustic phonon mode at the M point [see Fig.4(f)] and are strongest when the energy of this mode is close to zero, while the electron-phonon coupling is the strongest (see Supplementary Figure 2 in SI). When the temperature is elevated, the M-point acoustic phonon hardens, which quenches the electron-phonon scatterings between Se 4p and Ti 3d states. This further supports the correlation discussed above between the melting and recovery dynamics of the CDW fluctuating Se 4p band and hardening of the M phonon mode Otto et al. (2021). Similar microscopic explanation, which includes an active soft bosonic mode, was used many years ago to explain the fluctuating (or so-called pseudogap) region above the CDW phase Sadovskii (2001) in the quasi-1D Peierls systems Lee et al. (1973); McKenzie (1995), as well as in the high- cuprates Pines (1997), and more recently in metallic 2H-NbSe2 and 2H-TaSe2 Kuchinskii et al. (2012). Further, spectral weight redistribution around the Fermi level due to electron-phonon-related fluctuations for was once discussed for 1T-TiSe2 Yoshiyama et al. (1986), but it was never thoroughly appreciated and later connected to the replicas and other photoemission signatures in 1T-TiSe2. Our microscopic interpretation is also in line with the observation that the CDW fluctuations reach a maximum at , as observed with X-ray thermal diffuse scattering Holt et al. (2001). The temperature dependence of the phonon scattering lifetime characterized by hardening of the soft phonon mode with increasing temperature and a concomitant reduction of the electron scattering rates (i.e., ), consistent with transport measurements Velebit et al. (2016), is shown in Supplementary Figures 2 and 3. It shows the proportionality between the larger CDW gap and the longer lifetime of the amplitude mode. This is most likely caused by less quasiparticle damping and suggests that even in the short-range ordered CDW regions, there is a finite CDW gap that suppresses the damping of the coherent (and re-hardened) amplitude mode oscillation. Notably, the figures also show that artificial introduction of small periodic lattice distortions leads to an increase in the amplitude mode lifetime and better lifetime agreement with the experiment. However, the dynamical electron-phonon scattering alone redistributes the density of states enough to account for a pseudogap formation, as shown in Supplementary Figure 3 in SI. We therefore conclude that while static CDW gaps vanish above the transition temperature, the dynamic pseudogap persists. This affects both the electronic scattering and phonon suppression rates.
@@
Figure 4(f) additionally shows the phonon spectral function of 1T-TiSe2 in the fluctuation regime with included phonon-electron coupling, where part of the M phonon spectral weight extends up to negative frequencies. This indicates that in the fluctuation region where , the atomic motions fluctuate between normal and CDW states. Note that equilibrium electron diffraction measurements Cheng et al. (2022) as well as X-ray diffuse scattering Woo et al. (1976b); Guo et al. (2025) observed both L and M diffusive peaks at RT, suggesting the presence of the fluctuating periodic lattice distortions above . In Supplementary Figure 4 we demonstrate how the dynamical electron-phonon scattering introduces very similar modification to the electronic structure as the symmetry-breaking periodic lattice distortions in the CDW state. The system could, therefore, support the excitation of the CDW amplitude mode at , as it is observed experimentally following optical photoexcitation [Fig. 3(f)]. These theoretical results demonstrate that the features (i)-(iii) as well as the CDW amplitude mode observed in the CDW fluctuation regime could be explained by the dynamical electron-phonon effects (also supported by our scanning tunneling microscopy measurements - see Supplementary Figure 5 in SI), which opposes the common belief that the fluctuation regime is dominated by the electron-hole correlations Monney et al. (2012, 2015). The latter mechanism could still be present in the formation process of the CDW fluctuations in 1T-TiSe2, but more likely as a higher-order correcting effect. This is further corroborated by the fact that the equivalent electron-related collective soft mode (e.g., plasmon or exciton), which could produce the above features in scattering with Ti 3d or Se 4p states, is not observed in the normal phase above , neither in experiments Lin et al. (2022); Kogar et al. (2017) nor theoretical studies Lian et al. (2019).
@@
In conclusion, our study provides direct evidence of charge density wave (CDW) fluctuations in 1T-TiSe2 at room temperature using time-resolved extreme ultraviolet (XUV) momentum microscopy. We observe the persistence of the backfolded Se 4p band well above , demonstrating that CDW fluctuations remain detectable even at elevated temperatures. The ultrafast melting and recovery of these fluctuations, featuring the presence of coherent CDW amplitude phonon modes, highlight the importance of electron-phonon interactions in stabilizing the CDW phase at elevated temperatures.
@@
These results not only refine the understanding of CDW physics in 1T-TiSe2 but also have broader consequences for the study of fluctuating phases in other transition metal dichalcogenides and strongly correlated materials. Given the emerging connection between CDW fluctuations and unconventional superconductivity, our findings may provide valuable insights into the interplay between these two phenomena. Future studies leveraging advanced ultrafast techniques could further elucidate the microscopic mechanisms governing the fluctuating CDW phase, potentially unveiling new pathways for manipulating correlated electron systems in quantum materials.
@@ Methods
Experimental Methods. The experimental apparatus is centered around a custom-designed, polarization-tunable ultrafast XUV beamline Comby et al. (2022), interfaced with a Momentum Microscope end-station Medjanik et al. (2017). This beamline is driven by a commercial high-repetition-rate Yb fiber laser (1030 nm, 135 fs FWHM, 50 W) from Amplitude Laser Group, operating at 166 kHz. In the probe branch, a segment of the laser output is frequency-doubled using a BBO crystal, yielding 515 nm pulses (5 W) that are tightly focused into a thin, high-density argon gas jet configured in an annular geometry to facilitate high-order harmonic generation (HHG). Spatial filtering through pinholes effectively separates the intense 515 nm fundamental from the emerging XUV beamlet. To extract the 9th harmonic (21.6 eV), a combination of SiC/Mg multilayer mirrors (NTTAT) and a 200 nm Sn metallic filter is used for spectral selection. On the pump side, a small portion of the original 1030 nm laser pulse (135 fs FWHM) is directed as the IR excitation. The IR pump and XUV probe pulses are recombined collinearly via a drilled mirror and focused onto the target. Bulk iodine vapor transport-grown - single crystals Watson et al. (2019) are freshly cleaved and positioned on a motorized 6-axis hexapod for high-precision alignment within the main vacuum chamber, which maintains an ultra-high base pressure of 210-10 mbar. Photoemission data acquisition is carried out with a custom-designed time-of-flight momentum microscope featuring an advanced front lens capable of multiple operating modes (GST mbH) Tkach et al. (2024b). This detection system enables simultaneous mapping of the entire surface Brillouin zone over a broad binding energy range without necessitating sample orientation adjustments Medjanik et al. (2017). Measurements in this study were conducted at room temperature, where the total energy resolution reaches . This resolution, , results from the convolution of the Fermi-Dirac distribution width at a given temperature , the bandwidth of the ionizing radiation , and the time-of-flight detector resolution . The thermal broadening of the Fermi-Dirac distribution is estimated as , corresponding to approximately at room temperature. For data post-processing, an open-source workflow Xian et al. (2020, 2019) is employed to efficiently convert raw single-event-based datasets into calibrated, binned data hypervolumes, incorporating appropriate axis calibration. Further details about the experimental setup are provided in Ref. Fragkos et al. (2025).
@@
The native defects observed with STM on the sample perfectly match those observed in Arpaia et al. (2019), with an estimated defect density of . Only very few defects of type D associated to intercalated Ti atoms could be observed (density ), validating the low Ti doping of the sample. No signs of CDW modulation have been observed with STM even close to [Supplementary Figures 5(a) and 5(b) in SI]. Indeed, the FFT-amplitude plot [Supplementary Figure 5(c)] only shows spots corresponding to the unit cell charge modulation.
Computational Methods. Ground state calculations were performed within the framework of the DFT using the Quantum Espresso package Giannozzi et al. (2017). We use optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) with the PBE exchange-correlation functional Perdew et al. (1996) from the PseudoDojo library van Setten et al. (2018). The plane wave energy cutoff was fixed to Ry. To reduce the interaction between the periodic images we modeled the 2D layers in periodic cells with Å vacuum in the direction normal to the layer. The convergence criterion for energy is set to 10-8 eV and the atomic positions are relaxed until the Hellmann-Feynman forces are less than 10-4 eV/Å. A lattice parameter of is obtained with the PBE functional. The relaxed atomic positions in fractional (crystal) coordinates are Ti (0.0000000000, 0.0000000000, 0.5000000000), Se (0.3333333333, 0.6666666667, 0.4163666236), and Se (0.6666666667, 0.3333333333, 0.5836333764). To obtain the ground state charge density the 2D Brillouin zone is sampled with k-point mesh. In order to determine the , i.e., the temperature at which the relevant acoustic phonon mode becomes unstable, we use the Fermi-Dirac smearing functions with the electronic temperatures ranging from 100 K to 1600 K, while for the calculations of the electron self-energy we range from 1100 K to 1600 K.
@@
Data Availability
The ARPES and DFT data that support the findings of this study are available in Zenodo under the identifier https://doi.org/10.5281/zenodo.17813602. Additional data can be obtained from the authors upon reasonable request.